If -16 is added to a number, the sum will be 8 times the number. Find the number?
Twenty Five less than twice the number is -57.

Found 2 solutions by josgarithmetic, ankor@dixie-net.com:
Answer by josgarithmetic(39617)

--------------------------------------------------------
If -16 is added to a number, the sum will be 8 times the number. Find the number?
--------------------------------------------------------

n-16=8n



----------------------------------------------------
Twenty Five less than twice the number is -57.
----------------------------------------------------

2n-25=-57




-
Each was a different description. n used in both but not intended to be the same value for both. Literal translation into equations shown; but you solve each on your own.

Answer by ankor@dixie-net.com(22740)

If -16 is added to a number, the sum will be 8 times the number
-16 + x = 8x
Find the number?
-16 = 8x - x
-16 = 7x
x = -16/7 is the number
:
Check this solution in the original equation
Use the decimal equiv of -16/7; -2.2857
-16 -2.2857 = 8(-2.2857)
-18.2857 = -18.2857
:
;
Twenty Five less than twice the number is -57.
2x - 25 = -57
2x = -57 + 25
2x = -32
x = -32/2
x = -16